# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ Appendix B Mean-wave kinetic theory
$\Delta_{\bm{p}}^{s\tilde{s}}(t)\equiv\frac{1-e^{-i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}}{i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}.$ (31)
Here we will generally consider interactions between waves of opposite chiralities due to their coupling with the condensate mode $\bm{k}$, assuming $\langle a_{\bm{p}}^{s}a_{\bm{q}}^{-s}\rangle\neq 0$, and follow the same procedure as in [gome2025waves]. Note that such correlations are absent initially, as our forcing does not directly correlate waves of opposite polarity, $\langle f_{\bm{p}}^{s}f_{\bm{q}}^{-s}\rangle=0$). An additional equation is therefore needed to close the system, given by the evolution of the correlation between opposite polarity modes $\langle a_{\bm{p}}^{s*}a_{\bm{p}}^{-s}\rangle$
$\frac{\langle a_{\bm{p}}^{s*}a_{\bm{p}}^{-s}\rangle(t)-\langle a_{\bm{p}}^{s*}a_{\bm{p}}^{-s}\rangle(0)}{t}$ $=\sum_{\bm{q},\bm{k},s}\Bigg[U_{\bm{k}}^{*}V_{\bm{p}\bm{k}\bm{q}}^{-s,\tilde{s}}\langle a_{\bm{p}}^{s}a_{\bm{q}}^{\tilde{s}}\rangle^{*}\Delta^{-s,\tilde{s}}+U_{\bm{k}}V_{\bm{p}\bm{k}\bm{q}}^{s\tilde{s}*}\langle a_{\bm{p}}^{-s}a_{\bm{q}}^{\tilde{s}}\rangle\Delta^{s\tilde{s}*}\Bigg]\delta_{\bm{k}\bm{p}\bm{q}}$ (32)
@@
Similarly to [gome2025waves], we approximate the oscillating factor in equations (30) and (41) as
$\Delta_{\bm{p}}^{s\tilde{s}}(t)=\frac{1-e^{-i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}}{i\omega_{\bm{p}\bm{q}}^{s\tilde{s}}t}\approx F_{\bm{p}\bm{q}}\left(\frac{{U^{\prime}}}{\Omega}\right),$ (33)
@@
$F_{\bm{p}\bm{q}}$ $=\Theta\Big(1-\frac{|\omega_{\bm{p}\bm{q}}^{s\tilde{s}}|}{{U^{\prime}}}\Big)$ (34)
@@
$\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{s}<{U^{\prime}}$ (35)
@@
For cross-helicity interactions $\tilde{s}=-s$, using scale separation between $k$ and $p$ and recalling that $\bm{q}=-\bm{p}-\bm{k}$ we get $\omega_{\bm{p}\bm{q}}^{s,-s}\simeq 2\omega_{\bm{p}}^{s}=4s\Omega p_{z}/p$. Thus, the step-function approximation (33) results in a condition delimiting near-resonant from off-resonant interactions:
$\Delta_{\bm{p}}^{s,-s}(t)=\frac{1-e^{i\omega_{\bm{p}\bm{q}}^{s,-s}t}}{i\omega_{\bm{p}\bm{q}}^{s,-s}t}\approx\begin{cases}1,~~\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{-s}\leq{U^{\prime}}\Leftrightarrow~~\frac{4p_{z}}{p}\leq\frac{U^{\prime}}{\Omega}\\ 0,~~\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{-s}>{U^{\prime}}\Leftrightarrow~~\frac{4p_{z}}{p}>\frac{U^{\prime}}{\Omega}\end{cases}$ (36)
@@
For modes $\bm{p}$ in sector H, heterochiral-wave interactions decorrelate ($\Delta_{\bm{p}}^{s,-s}\approx 0$ ) and only homochiral-wave interactions ($\tilde{s}=s$) are present. For modes in sector A, on the other hand, both types of interactions are allowed by the resonance condition and these modes experience the full 3D QL dynamics. Note that the resonant condition for homochiral-wave interactions (35) is less restrictive than that for heterochiral-wave interactions (36), and that the former becomes restrictive for modes in sector H at sufficiently high rotation rates. At large rotation ($\frac{{U^{\prime}}}{\Omega}<l_{f}/(L_{y})$), modes for which $\Delta_{\bm{p}}^{ss}\approx 0$ at the forcing scale do not interact with the condensate, as discussed in [gome2025waves], and in the QL dynamics are treated as if unforced.
Note also that there is a lower bound for the time $t$ we need to consider here: the oscillations of the waves will be felt by the system (meaning that some interactions will average out to zero) only for times much larger than $1/\max(\omega_{\bm{p}\bm{q}}^{s\tilde{s}})=1/4\Omega$, which occurs in interactions between opposite-helicity waves, $s=-\tilde{s}$. Hence, such an expansion makes sense for $1/4\Omega\ll t\ll 1/U^{\prime}$, requiring a sufficient time-scale separation. Thus, the kinetic equation that we derive is an expansion in $U^{\prime}/4\Omega$ (the ratio between the wave and non-linear time scale), where we keep the leading order term $O(U^{\prime}/4\Omega)$. Condition $U^{\prime}/4\Omega<1$ is required for the existence of Sector H.
We now consider the lowest complex-conjugate modes $k_{y}=\pm 2\pi/L_{y}$ for the condensate $U_{\bm{k}}=\frac{{U^{\prime}}}{i\sqrt{2}k_{y}}$, and expand the coefficients in the limit $Lk_{f}\gg 1$:
$V_{\bm{p}\bm{k}\bm{q}}^{ss}$ $=-ip_{x}-sk_{y}\left(\frac{p_{x}^{2}p_{z}}{p_{\perp}^{2}p}+\frac{p_{z}}{2p}\right)-ik_{y}\frac{p_{x}p_{y}}{2p^{2}}+O\left(\frac{k_{y}^{2}}{p^{2}}\right),~~~~~~~~~~~V_{\bm{p}\bm{k}\bm{q}}^{s,-s}=ik_{y}V_{\bm{p}}^{s,-s}+O\left(\frac{k_{y}^{2}}{p^{2}}\right),$ (38)
$=-ip_{x}-2ik_{y}V_{\bm{p}}^{ss}+O\left(\frac{k_{y}^{2}}{p^{2}}\right),$
@@
$V_{\bm{p}}^{ss}\equiv-is\frac{p_{z}}{p}\left(1+\frac{2p_{x}^{2}}{p_{\perp}^{2}}\right)+\frac{p_{x}p_{y}}{p^{2}},~~~~~~~~~~~V_{\bm{p}}^{s,-s}\equiv\frac{p_{x}p_{y}}{p^{2}}\Big(1+\frac{p_{z}^{2}}{p_{\perp}^{2}}\Big)+is\frac{p_{z}}{p}\Big(1-\frac{2p_{x}^{2}}{p_{\perp}^{2}}\Big).$ (39)
When considering condition (36) for heterochiral-wave interactions, and assuming that homochiral waves resonate with the condensate ($\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{s}<{U^{\prime}}$), we obtain the coupled equations
$\partial_{t}e_{\bm{p}}^{s}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\Big[p_{x}\partial_{p_{y}}+\Re[V_{\bm{p}}^{ss}]\Big]\Phi_{\bm{p}\bm{k}}^{s}-\mathbb{1}_{\bm{p}\in A}~\Re(V_{\bm{p}}^{s,-s}\Psi_{\bm{p}\bm{k}}^{s,-s*})\Bigg)+\epsilon\chi_{\bm{p}}^{s}$ (40)
$\partial_{t}\langle a_{\bm{p}}^{s*}a_{\bm{p}}^{-s}\rangle$ $=2\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\Big[p_{x}\partial_{p_{y}}+V_{\bm{p}}^{ss}\Big](\Psi_{\bm{p}\bm{k}}^{s,-s*}+\Psi_{\bm{p}\bm{k}}^{-s,s})~~~~-\mathbb{1}_{\bm{p}\in A}~V_{\bm{p}}^{s,-s*}(\Phi_{\bm{p}\bm{k}}^{s*}+\Phi_{\bm{p}\bm{k}}^{-s})\Bigg),$ (41)
@@
$\begin{split}\Phi_{\bm{p}\bm{k}}^{s}&\equiv\frac{\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{s}\rangle}{2}=\frac{\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle+c.c}{2}+O\left(\frac{k_{y}}{p}\right)\end{split}$ (42)
@@
$\begin{split}\Psi_{\bm{p}\bm{k}}^{s,-s}\equiv\frac{\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{-s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{-s}\rangle}{2}\end{split}.$ (43)
Note that $\Im(\Phi_{\bm{p}\bm{k}})=O(k_{y}/p)$, while $\Psi_{\bm{p}\bm{k}}\in\mathbb{C}$ at the leading order in $k_{y}/p$.
The wave kinetic equations (40)-(41) are coupled with the mean-flow equation (26), obtained by Reynolds averaging over time $T\gg 1/{U^{\prime}}$ and over $z$, in a time window where the ensemble-averaged correlator $\langle a_{\bm{p}}^{s}a_{\bm{q}}^{s^{\prime}}\rangle$ is stationary. The total energy transfer is
$T_{\rm 3D-2D}$ $=\sum_{\bm{p},s,\tilde{s}}U^{\prime}_{-\bm{k}}\langle u_{\bm{p}}^{s}v_{-\bm{p}+\bm{k}}^{\tilde{s}}\rangle+U^{\prime}_{\bm{k}}\langle u_{\bm{p}}^{s}v_{-\bm{p}-\bm{k}}^{\tilde{s}}\rangle=\frac{\mathcal{V}{U^{\prime}}}{\sqrt{2}}\int{\rm d}\bm{p}\sum_{s,\tilde{s}=\pm 1}\langle uv\rangle^{s\tilde{s}}_{\bm{p}\bm{k}}$ (44)
@@
$\langle uv\rangle_{\bm{p}\bm{k}}^{ss}\equiv\frac{1}{2}\Big(\langle u_{\bm{p}}^{s}v^{s}_{-\bm{p}-\bm{k}}\rangle+\langle u_{\bm{p}}^{s}v_{-\bm{p}+\bm{k}}^{s}\rangle\Big)+c.c=-\frac{p_{x}p_{y}}{p^{2}}~\Phi_{\bm{p}\bm{k}}^{s}~+O\left(\frac{k_{y}}{p}\right)$ (45)
where in the last line we used that $h_{\bm{p},x}^{-s}h_{-\bm{p}-\bm{k},y}^{-s}=h_{\bm{p},x}^{-s}h_{-\bm{p},y}^{-s}+O(k_{y}/p)=-\frac{p_{x}p_{y}}{2p^{2}}+is\frac{p_{z}}{2p}+O(k_{y}/p)$, that $\Phi_{\bm{p}\bm{k}}^{s}=\Phi_{\bm{p},-\bm{k}}^{s}$ and that the imaginary part of the correlator is subleading:
$2\Im\left[\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{s}\rangle\right]$ $=\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{s}\rangle-c.c$
$=\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle-\langle a_{\bm{p}+\bm{k}}^{s}a_{-\bm{p}}^{s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{s}\rangle-\langle a_{\bm{p}-\bm{k}}^{s}a_{-\bm{p}}^{s}\rangle$
$=-k_{y}\partial_{p_{y}}\langle a_{\bm{p}-\bm{k}}^{s}a_{-\bm{p}}^{s}\rangle+k_{y}\partial_{p_{y}}\langle a_{\bm{p}-\bm{k}}^{s}a_{-\bm{p}}^{s}\rangle$
@@
$\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}$ $\equiv\langle u_{\bm{p}}^{s}v_{-\bm{p}+\bm{k}}^{-s}+u_{\bm{p}}^{s}v_{-\bm{p}-\bm{k}}^{-s}\rangle$
$=\frac{1}{T}\int_{0}^{T}~dt~\Big(h_{\bm{p},x}^{-s}h_{-\bm{p}+\bm{k},y}^{s}\langle a_{\bm{p}}^{-s*}a_{-\bm{p}+\bm{k}}^{s*}\rangle e^{i(\omega_{\bm{p}}^{s}+\omega_{-\bm{p}+\bm{k}}^{-s})t}+(\bm{k}\to-\bm{k})\Big)$
$\simeq\Re(V_{\bm{p}}^{s,-s}\Psi_{\bm{p}\bm{k}}^{s,-s})+O\left(\frac{k_{y}}{p}\right)$ (47)
@@
$\partial_{t}e_{\bm{p}}^{s}+\partial_{p_{y}}\Pi^{s}_{\bm{p}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\left(\langle uv\rangle_{\bm{p}\bm{k}}^{ss}+\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}\right)+\epsilon\chi_{\bm{p}}^{s}$ (48)
@@
$\partial_{p_{y}}\Pi^{s}_{\bm{p}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}+\epsilon\chi_{\bm{p}}^{s},$ (49)
with the total Reynolds stress correlator $\langle uv\rangle_{\bm{p}\bm{k}}=\langle uv\rangle_{\bm{p}\bm{k}}^{ss}+\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}$. Away from the forcing scale ($\chi_{\bm{p}}^{s}=0$), waves either lose energy to the 2D flow (the wave energy flux $\Pi_{\bm{k}}^{s}$ decreases with $p_{y}$ and the transfer to the 2D flow $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}>0$), or gain energy from the 2D flow.

### v2

@@ Appendix B Mean-wave kinetic theory
$\Delta_{\boldsymbol{p}}^{s\tilde{s}}(t)\equiv\frac{1-e^{-i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}}{i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}.$ (31)
Here we will generally consider interactions between waves of opposite chiralities due to their coupling with the condensate mode $\boldsymbol{k}$, assuming $\langle a_{\boldsymbol{p}}^{s}a_{\boldsymbol{q}}^{-s}\rangle\neq 0$, and follow the same procedure as in [26]. Note that such correlations are absent initially, as our forcing does not directly correlate waves of opposite polarity, $\langle f_{\boldsymbol{p}}^{s}f_{\boldsymbol{q}}^{-s}\rangle=0$). An additional equation is therefore needed to close the system, given by the evolution of the correlation between opposite polarity modes $\langle a_{\boldsymbol{p}}^{s*}a_{\boldsymbol{p}}^{-s}\rangle$
$\frac{\langle a_{\boldsymbol{p}}^{s*}a_{\boldsymbol{p}}^{-s}\rangle(t)-\langle a_{\boldsymbol{p}}^{s*}a_{\boldsymbol{p}}^{-s}\rangle(0)}{t}$ $=\sum_{\boldsymbol{q},\boldsymbol{k},s}\Bigg[U_{\boldsymbol{k}}^{*}V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{-s,\tilde{s}}\langle a_{\boldsymbol{p}}^{s}a_{\boldsymbol{q}}^{\tilde{s}}\rangle^{*}\Delta^{-s,\tilde{s}}+U_{\boldsymbol{k}}V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s\tilde{s}*}\langle a_{\boldsymbol{p}}^{-s}a_{\boldsymbol{q}}^{\tilde{s}}\rangle\Delta^{s\tilde{s}*}\Bigg]\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}$ (32)
@@
Similarly to [26], we approximate the oscillating factor in equations (30) and (41) as
$\Delta_{\boldsymbol{p}}^{s\tilde{s}}(t)=\frac{1-e^{-i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}}{i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}t}\approx F_{\boldsymbol{p}\boldsymbol{q}}\left(\frac{{U^{\prime}}}{\Omega}\right),$ (33)
@@
$F_{\boldsymbol{p}\boldsymbol{q}}$ $=\Theta\Big(1-\frac{|\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}}|}{{U^{\prime}}}\Big)$ (34)
@@
$\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{s}<{U^{\prime}}$ (35)
@@
For cross-helicity interactions $\tilde{s}=-s$, using scale separation between $k$ and $p$ and recalling that $\boldsymbol{q}=-\boldsymbol{p}-\boldsymbol{k}$ we get $\omega_{\boldsymbol{p}\boldsymbol{q}}^{s,-s}\simeq 2\omega_{\boldsymbol{p}}^{s}=4s\Omega p_{z}/p$. Thus, the step-function approximation (33) results in a condition delimiting near-resonant from off-resonant interactions:
$\Delta_{\boldsymbol{p}}^{s,-s}(t)=\frac{1-e^{i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s,-s}t}}{i\omega_{\boldsymbol{p}\boldsymbol{q}}^{s,-s}t}\approx\begin{cases}1,~~\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{-s}\leq{U^{\prime}}\Leftrightarrow~~\frac{4p_{z}}{p}\leq\frac{U^{\prime}}{\Omega}\\ 0,~~\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{-s}>{U^{\prime}}\Leftrightarrow~~\frac{4p_{z}}{p}>\frac{U^{\prime}}{\Omega}\end{cases}$ (36)
@@
For modes $\boldsymbol{p}$ in sector H, heterochiral-wave interactions decorrelate ($\Delta_{\boldsymbol{p}}^{s,-s}\approx 0$ ) and only homochiral-wave interactions ($\tilde{s}=s$) are present. For modes in sector A, on the other hand, both types of interactions are allowed by the resonance condition and these modes experience the full 3D QL dynamics. Note that the resonant condition for homochiral-wave interactions (35) is less restrictive than that for heterochiral-wave interactions (36), and that the former becomes restrictive for modes in sector H at sufficiently high rotation rates. At large rotation ($\frac{{U^{\prime}}}{\Omega}<l_{f}/(L_{y})$), modes for which $\Delta_{\boldsymbol{p}}^{ss}\approx 0$ at the forcing scale do not interact with the condensate, as discussed in [26], and in the QL dynamics are treated as if unforced.
Note also that there is a lower bound for the time $t$ we need to consider here: the oscillations of the waves will be felt by the system (meaning that some interactions will average out to zero) only for times much larger than $1/\max(\omega_{\boldsymbol{p}\boldsymbol{q}}^{s\tilde{s}})=1/4\Omega$, which occurs in interactions between opposite-helicity waves, $s=-\tilde{s}$. Hence, such an expansion makes sense for $1/4\Omega\ll t\ll 1/U^{\prime}$, requiring a sufficient time-scale separation. Thus, the kinetic equation that we derive is an expansion in $U^{\prime}/4\Omega$ (the ratio between the wave and non-linear time scale), where we keep the leading order term $O(U^{\prime}/4\Omega)$. Condition $U^{\prime}/4\Omega<1$ is required for the existence of Sector H.
We now consider the lowest complex-conjugate modes $k_{y}=\pm 2\pi/L_{y}$ for the condensate $U_{\boldsymbol{k}}=\frac{{U^{\prime}}}{i\sqrt{2}k_{y}}$, and expand the coefficients in the limit $Lk_{f}\gg 1$:
$V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{ss}$ $=-ip_{x}-sk_{y}\left(\frac{p_{x}^{2}p_{z}}{p_{\perp}^{2}p}+\frac{p_{z}}{2p}\right)-ik_{y}\frac{p_{x}p_{y}}{2p^{2}}+O\left(\frac{k_{y}^{2}}{p^{2}}\right),~~~~~~~~~~~V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s,-s}=ik_{y}V_{\boldsymbol{p}}^{s,-s}+O\left(\frac{k_{y}^{2}}{p^{2}}\right),$ (38)
$=-ip_{x}-2ik_{y}V_{\boldsymbol{p}}^{ss}+O\left(\frac{k_{y}^{2}}{p^{2}}\right),$
@@
$V_{\boldsymbol{p}}^{ss}\equiv-is\frac{p_{z}}{p}\left(1+\frac{2p_{x}^{2}}{p_{\perp}^{2}}\right)+\frac{p_{x}p_{y}}{p^{2}},~~~~~~~~~~~V_{\boldsymbol{p}}^{s,-s}\equiv\frac{p_{x}p_{y}}{p^{2}}\Big(1+\frac{p_{z}^{2}}{p_{\perp}^{2}}\Big)+is\frac{p_{z}}{p}\Big(1-\frac{2p_{x}^{2}}{p_{\perp}^{2}}\Big).$ (39)
When considering condition (36) for heterochiral-wave interactions, and assuming that homochiral waves resonate with the condensate ($\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{s}<{U^{\prime}}$), we obtain the coupled equations
$\partial_{t}e_{\boldsymbol{p}}^{s}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\Big[p_{x}\partial_{p_{y}}+\Re[V_{\boldsymbol{p}}^{ss}]\Big]\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}-\mathbb{1}_{\boldsymbol{p}\in A}~\Re(V_{\boldsymbol{p}}^{s,-s}\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s,-s*})\Bigg)+\epsilon\chi_{\boldsymbol{p}}^{s}$ (40)
$\partial_{t}\langle a_{\boldsymbol{p}}^{s*}a_{\boldsymbol{p}}^{-s}\rangle$ $=2\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\Big[p_{x}\partial_{p_{y}}+V_{\boldsymbol{p}}^{ss}\Big](\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s,-s*}+\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-s,s})~~~~-\mathbb{1}_{\boldsymbol{p}\in A}~V_{\boldsymbol{p}}^{s,-s*}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s*}+\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-s})\Bigg),$ (41)
@@
$\begin{split}\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}&\equiv\frac{\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{s}\rangle}{2}=\frac{\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle+c.c}{2}+O\left(\frac{k_{y}}{p}\right)\end{split}$ (42)
@@
$\begin{split}\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}\equiv\frac{\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{-s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{-s}\rangle}{2}\end{split}.$ (43)
Note that $\Im(\Phi_{\boldsymbol{p}\boldsymbol{k}})=O(k_{y}/p)$, while $\Psi_{\boldsymbol{p}\boldsymbol{k}}\in\mathbb{C}$ at the leading order in $k_{y}/p$.
The wave kinetic equations (40)-(41) are coupled with the mean-flow equation (26), obtained by Reynolds averaging over time $T\gg 1/{U^{\prime}}$ and over $z$, in a time window where the ensemble-averaged correlator $\langle a_{\boldsymbol{p}}^{s}a_{\boldsymbol{q}}^{s^{\prime}}\rangle$ is stationary. The total energy transfer is
$T_{\rm 3D-2D}$ $=\sum_{\boldsymbol{p},s,\tilde{s}}U^{\prime}_{-\boldsymbol{k}}\langle u_{\boldsymbol{p}}^{s}v_{-\boldsymbol{p}+\boldsymbol{k}}^{\tilde{s}}\rangle+U^{\prime}_{\boldsymbol{k}}\langle u_{\boldsymbol{p}}^{s}v_{-\boldsymbol{p}-\boldsymbol{k}}^{\tilde{s}}\rangle=\frac{\mathcal{V}{U^{\prime}}}{\sqrt{2}}\int{\rm d}\boldsymbol{p}\sum_{s,\tilde{s}=\pm 1}\langle uv\rangle^{s\tilde{s}}_{\boldsymbol{p}\boldsymbol{k}}$ (44)
@@
$\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{ss}\equiv\frac{1}{2}\Big(\langle u_{\boldsymbol{p}}^{s}v^{s}_{-\boldsymbol{p}-\boldsymbol{k}}\rangle+\langle u_{\boldsymbol{p}}^{s}v_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle\Big)+c.c=-\frac{p_{x}p_{y}}{p^{2}}~\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}~+O\left(\frac{k_{y}}{p}\right)$ (45)
where in the last line we used that $h_{\boldsymbol{p},x}^{-s}h_{-\boldsymbol{p}-\boldsymbol{k},y}^{-s}=h_{\boldsymbol{p},x}^{-s}h_{-\boldsymbol{p},y}^{-s}+O(k_{y}/p)=-\frac{p_{x}p_{y}}{2p^{2}}+is\frac{p_{z}}{2p}+O(k_{y}/p)$, that $\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=\Phi_{\boldsymbol{p},-\boldsymbol{k}}^{s}$ and that the imaginary part of the correlator is subleading:
$2\Im\left[\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{s}\rangle\right]$ $=\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{s}\rangle-c.c$
$=\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle-\langle a_{\boldsymbol{p}+\boldsymbol{k}}^{s}a_{-\boldsymbol{p}}^{s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{s}\rangle-\langle a_{\boldsymbol{p}-\boldsymbol{k}}^{s}a_{-\boldsymbol{p}}^{s}\rangle$
$=-k_{y}\partial_{p_{y}}\langle a_{\boldsymbol{p}-\boldsymbol{k}}^{s}a_{-\boldsymbol{p}}^{s}\rangle+k_{y}\partial_{p_{y}}\langle a_{\boldsymbol{p}-\boldsymbol{k}}^{s}a_{-\boldsymbol{p}}^{s}\rangle$
@@
$\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}$ $\equiv\langle u_{\boldsymbol{p}}^{s}v_{-\boldsymbol{p}+\boldsymbol{k}}^{-s}+u_{\boldsymbol{p}}^{s}v_{-\boldsymbol{p}-\boldsymbol{k}}^{-s}\rangle$
$=\frac{1}{T}\int_{0}^{T}~dt~\Big(h_{\boldsymbol{p},x}^{-s}h_{-\boldsymbol{p}+\boldsymbol{k},y}^{s}\langle a_{\boldsymbol{p}}^{-s*}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s*}\rangle e^{i(\omega_{\boldsymbol{p}}^{s}+\omega_{-\boldsymbol{p}+\boldsymbol{k}}^{-s})t}+(\boldsymbol{k}\to-\boldsymbol{k})\Big)$
$\simeq\Re(V_{\boldsymbol{p}}^{s,-s}\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s,-s})+O\left(\frac{k_{y}}{p}\right)$ (47)
@@
$\partial_{t}e_{\boldsymbol{p}}^{s}+\partial_{p_{y}}\Pi^{s}_{\boldsymbol{p}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\left(\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{ss}+\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}\right)+\epsilon\chi_{\boldsymbol{p}}^{s}$ (48)
@@
$\partial_{p_{y}}\Pi^{s}_{\boldsymbol{p}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}+\epsilon\chi_{\boldsymbol{p}}^{s},$ (49)
with the total Reynolds stress correlator $\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}=\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{ss}+\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}$. Away from the forcing scale ($\chi_{\boldsymbol{p}}^{s}=0$), waves either lose energy to the 2D flow (the wave energy flux $\Pi_{\boldsymbol{k}}^{s}$ decreases with $p_{y}$ and the transfer to the 2D flow $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}>0$), or gain energy from the 2D flow.